# Can Graph Learning Improve Task Planning?

### v1

@@ Abstract
Task planning is emerging as an important research topic alongside the development of large language models (LLMs). It aims to break down complex user requests into solvable sub-tasks, thereby fulfilling the original requests. In this context, the sub-tasks can be naturally viewed as a graph, where the nodes represent the sub-tasks, and the edges denote the dependencies among them. Consequently, task planning is a decision-making problem that involves selecting a connected path or subgraph within the corresponding graph and invoking it. In this paper, we explore graph learning-based methods for task planning, a direction that is orthogonal to the prevalent focus on prompt design. Our interest in graph learning stems from a theoretical discovery: the biases of attention and auto-regressive loss impede LLMs’ ability to effectively navigate decision-making on graphs, which is adeptly addressed by graph neural networks (GNNs). This theoretical insight led us to integrate GNNs with LLMs to enhance overall performance. Extensive experiments demonstrate that GNN-based methods surpass existing solutions even without training, and minimal training can further enhance their performance. Additionally, our approach complements prompt engineering and fine-tuning techniques, with performance further enhanced by improved prompts or a fine-tuned model 111 The code and datasets are available at https://github.com/WxxShirley/GNN4TaskPlan.
@@ 1 Introduction
LLM-empowered autonomous agents have recently emerged as a rapidly growing field of research and are considered a significant step towards artificial general intelligence (AGI) [43, 4]. These agents have achieved remarkable successes across a variety of domains, as evidenced by their ability to address complex AI challenges (e.g., HuggingGPT [31] and Visual ChatGPT [46]), excel in gaming environments (e.g., Voyager [40]), and drive innovation in chemical research (e.g., [3]). Within this burgeoning field, task planning emerges as a critical area of study. It involves LLMs autonomously interpreting user instructions, breaking these instructions down into concrete and solvable sub-tasks, and then fulfilling the user’s request by executing each sub-task [29, 31, 30]. For instance, in the case of HuggingGPT [31], task planning involves invoking expert AI models from the HuggingFace website to solve complex AI tasks beyond the capabilities of GPT alone.
@@
Given its practical significance, numerous algorithms have been proposed, with a major focus on prompt design [30, 31, 23, 20, 42, 34, 53, 2, 45]. This paper proposes to explore an orthogonal direction, i.e., graph-learning-based approaches. In task planning, solvable sub-tasks can be naturally represented as a task graph, wherein each node corresponds to a distinct sub-task, and each edge signifies the dependencies between these sub-tasks. The crux of task planning, therefore, involves selecting a connected path or subgraph to satisfy the user’s request, which is a decision-making problem on graphs. Adopting this framework, we analyze the task planning capabilities of LLMs, specifically within the context of HuggingGPT [31]. Our empirical investigation uncovers that a considerable portion of planning failures can be ascribed to the LLMs’ inefficacy in accurately discerning the structure of the task graph. This finding presents intriguing questions from both theoretical and empirical perspectives. Theoretically, it initiates a discussion on the inherent limitations of LLMs in processing graph-based tasks. Empirically, it highlights the urgent need for developing effective and efficient strategies to mitigate this deficiency and improve task planning performance.
@@
For the theoretical question, we first investigate the expressiveness of Transformer architectures when applied to graph tasks with sequential graph input, such as edge list representations. Our initial hypothesis is that the format of sequential graph input might not align with the inductive bias inherent to graph structures, potentially reducing expressiveness. Contrary to this hypothesis, it is proved that by taking edge lists as the input, a constant-width Transformer can solve graph decision-making problems by simulating dynamic programming algorithms on edge lists. Nevertheless, we find that LLMs’ solutions lack invariance under graph isomorphism, an important property for graph decision-making problems. In addition, the expressiveness is weakened if the attention is sparse, which is typically observed in LLMs [48]. Beyond expressiveness, we also examine the influence of auto-regressive loss, demonstrating that it introduces spurious correlations that can be harmful to graph decision-making tasks. These insights expose the inherent limitations of LLMs in task planning and, more broadly, in graph-related problems (e.g., the challenges in [10, 41, 24]).
@@
2 Preliminaries and Formulations
In this section, we introduce the task planning and GNNs. More importantly, we connect them with the classic concepts, i.e., decision-making problems and dynamic programming. This connection will play a pivotal role in the theoretical analysis and algorithm development.
2.1 Task Planning and Decision Making
In task planning, the input comprises an ambiguous user request along with a pool of available sub-tasks. The output is a selection of specific sub-tasks to be invoked and the manner in which they are to be executed. An illustrative example within HuggingGPT is depicted in Figure 1.
The sub-tasks form a task graph $G=(V,E)$ [30], where the nodes $V=\{v_{1},\cdots,v_{n}\}$ represent a set of solvable tasks, and an edge from $v_{i}$ to $v_{j}$ is established if the sub-task $v_{j}$ is dependent on the sub-task $v_{i}$. The goal of task planning is to identify a path or a connected sub-graph within the task graph that fulfills the users’ requests. In the context of HuggingGPT [31], the nodes correspond to expert AI models from the HuggingFace website. An edge is drawn from $v_{i}$ to $v_{j}$ if the input format of $v_{j}$ matches the output format of $v_{i}$. Task planning involves invoking a sequence of AI models (i.e., a path). Viewed from this angle, task planning bears resemblance to traditional decision-making problems on graphs, such as planning for the shortest path. The key distinction from classic decision-making problems lies in the specification of the objective value: it is explicitly defined in the former, while in task planning, it is implicitly inferred from the users’ requests.
@@ 2.1 Task Planning and Decision Making
Figure 1: Illustration of Task Planning in HuggingGPT.
2.2 GNNs and Dynamic Programming
GNNs are a class of neural networks preserving equivariance under graph isomorphism. The most popular GNNs are based on message passing scheme, where the node representation is iteratively updated by aggregating the representations from its neighbors. Formally, the $k$-th layer of a GNN is given by
$\bm{m}_{u}^{(k)}=\text{Aggregate}^{(k)}(\{[\bm{h}_{v}^{(k-1)},\bm{e}_{u,v}]|v\in\mathcal{N}(u)\}),\quad\bm{h}_{u}^{(k)}=\text{Combine}^{(k)}(\bm{h}_{u}^{(k-1)},\bm{m}^{(k)}_{u}),$ (1)
where $\mathcal{N}(u)$ denotes the set of in-neighbors of node $u$, $\bm{e}_{u,v}$ is the edge feature, and $\bm{h}_{u}^{(k)}$ is the representation of node $u$ with $\bm{h}_{u}^{(0)}$ representing initial node features. The aggregation function $\text{Aggregate}^{(k)}$ and update function $\text{Combine}^{(k)}$ are parameterized by neural networks and the choice of these two functions leads to different types of GNNs.
DP [1] presents a general framework for solving decision-making problems. It breaks a complex problem into simpler sub-tasks, each of which can be solved based on the results of previous sub-tasks. The general form of DP is expressed as
$\text{Answer}[k][i]=f\left(\square_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]\oplus c[i][j])\right),$ (2)
where $\text{Answer}[k][i]$ is the solution to state $i$ in the $k$-th iteration, $f$ is a task-specific update function, $c[i][j]$ is a cost associated with state $i$ and $j$, $\mathcal{T}(i)$ is the set of state can be transited to $i$, $\square_{j\in\mathcal{T}(i)}$ is an aggregation function such as MAX or SUM, $\oplus$ is an operator such as + or XOR. We give the formulation of typical DP algorithms in Appendix C.1. DP updates in (2) can be expressed by GNNs’ updates in Eq. (1) if we set the states as the nodes, $\mathcal{N}(u)=\mathcal{T}(u)$, and $\bm{e}_{u,v}=c[u][v]$. More formal relationships are established in [50, 7], showing that GNNs are well-suited for solving DP problems.
Figure 2: Illustration of performance and hallucination ratio across LLMs.
3 Diagnosing Task Planning in LLMs
3.1 Empirical Observations
@@
We adopt the experimental settings as outlined in the work of HuggingGPT [31], where the prompts are specifically optimized for task planning on HuggingFace. The evaluation metric calculates the F1 score to assess the accuracy of the tasks identified by LLMs against the ground-truth tasks. Additionally, we report two task-graph-related metrics: the node hallucination ratio and the edge hallucination ratio. These metrics measure the frequency of non-existent nodes (i.e., tasks) and edges (i.e., relations) outputted by LLMs, respectively, indicative of the models’ misinterpretation of the graph input.
@@
3.2 Theoretical Insights
Besides the example mentioned above, existing studies demonstrate that LLMs struggle with understanding graph-structured inputs [41, 10]. Despite these empirical evaluations, there remains a limited theoretical understanding of this phenomenon. In this subsection, we explore the performance of LLMs in graph decision-making problems, including task planning. In contrast to previous graph learning approaches for graph decision-making problems, LLMs process the graph input by flattening it into a sequence and are trained using an auto-regressive loss. We will then examine the impact of these two factors.
Sequential Graph Input: We consider general graph decision-making problems that can be resolved using DP as described in (2). The input comprises the edge list and initial states:
@@ 3.2 Theoretical Insights
The intended output format is $u_{1}\ \text{Answer}[k][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[k][u_{n}]$. To our surprise, although the edge list input does not directly reflect the geometric structures of graphs, it enables Transformers to simulate DP efficiently, in terms of expressiveness, as demonstrated by the following theorem.
@@
(Expressiveness) Assume the input format is given in (3) and $f,\oplus,\square$ in DP update (2) satisfy the assumptions 1 and 2. There exists a log-precision constant-depth and constant-width Transformer that simulates one step of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
The proof is presented in Appendix B.1. However, certain aspects of the proof’s constructions are challenging to be realized in Transformers that have been pretrained on natural language. First, the embedding process must be carefully filtered to ensure invariance under graph isomorphism. This invariance property does not align with the inductive biases inherent in natural language, making it difficult to achieve. Consequently, if an LLM can accurately produce the correct answer for a specific ordering of nodes, it might not maintain this accuracy after the nodes have been reordered (experiments given in Appendix C.2). Second, each token needs to synchronize its hidden states with all other tokens sharing the same token ID, which is of order $O(|V|)$. In practice, the attention trained from natural language is typically sparse [48], leading to intractability issues. The formal lower bound is provided in the following proposition and the proof is given in Appendix B.2.
@@
Assume the input format is described (3) and that the attention mechanism is limited to attending to a constant number of tokens. There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
Auto-regressive Loss: Our investigation next focuses on the auto-regressive loss and considers the following scenario: given a fixed task graph, user data is collected to perform instruction tuning with next-token-prediction loss. A question arises concerning whether auto-regressive loss might introduce spurious correlations harmful to planning problems. We conceptualize this issue through the lens of a path-planning problem. The training dataset comprises input sequences of the form $s\ t\ s\ v_{1}\ v_{2}\ \cdots\ t$, where $s$ represents the source node, $t$ the target node, and the sequence $s\ v_{1}\ v_{2}\ \cdots\ t$ is a path that adheres to specified constraints. During testing, given the initial and target nodes $s$ and $t$, the model is expected to generate a path with the same constraint. Our findings indicate that auto-regressive loss can lead to the emergence of a frequency-based spurious correlation, as substantiated by the following theorem and the proof is given in Appendix B.3.
@@
For a human, finding a path from $a$ to $d$ simply involves concatenating the paths $a\ b\ c$ and $b\ c\ d$. However, auto-regressive loss fails under such circumstances. In task planning datasets, we indeed observe that the performance of finetuned LLMs is inferior to that of GNNs trained on the same dataset, as shown in Figure 3(b).
4 A Simple Fix: Introducing GNNs
In contrast to LLMs, GNNs can strictly operate on the task graph, thereby avoiding hallucinations. Additionally, they leverage the graph structure as input, rather than flattening the graph into a sequence, thus overcoming the theoretical limitations discussed previously. Furthermore, GNNs have demonstrated proficiency in handling graph decision-making problems, both theoretically and empirically [50, 7, 16]. As a result, the simplest fix is to integrate GNNs into the task-planning algorithm. In this section, we first briefly introduce the current LLM-based solutions and then illustrate how GNNs can be integrated into them. We propose both training-free and training-based approaches to enhance performance. Training-free methods are necessary when the available tasks are continuously changing, or new tasks are emerging consistently. This scenario is common when the task planning module is deployed in a new system. Once the task planning module has been deployed for a period, it becomes possible to collect users’ requests and label a small proportion of the data, enabling lightweight training-based methods.
4.1 Purely LLM-based Task Planning
We outline a two-stage pipeline for task planning adopted in HuggingGPT [30, 31]. The first stage involves request decomposition, where a user’s ambiguous request is broken down into concrete steps via LLMs. For instance, the request illustrated in Figure 1 is decomposed into the following steps: (1) analyze the pose of the boy; (2) take that pose and generate a new image; (3) generate the caption for newly generated image; (4) convert the generated text into audio. The second stage is task retrieval. For each decomposed step, LLMs are employed to retrieve an appropriate task from the task pool and execute them in sequence.

### v2

# Can Graph Learning Improve Planning in LLM-based Agents?

@@ Abstract
Task planning in language agents is emerging as an important research topic alongside the development of large language models (LLMs). It aims to break down complex user requests in natural language into solvable sub-tasks, thereby fulfilling the original requests. In this context, the sub-tasks can be naturally viewed as a graph, where the nodes represent the sub-tasks, and the edges denote the dependencies among them. Consequently, task planning is a decision-making problem that involves selecting a connected path or subgraph within the corresponding graph and invoking it. In this paper, we explore graph learning-based methods for task planning, a direction that is orthogonal to the prevalent focus on prompt design. Our interest in graph learning stems from a theoretical discovery: the biases of attention and auto-regressive loss impede LLMs’ ability to effectively navigate decision-making on graphs, which is adeptly addressed by graph neural networks (GNNs). This theoretical insight led us to integrate GNNs with LLMs to enhance overall performance. Extensive experiments demonstrate that GNN-based methods surpass existing solutions even without training, and minimal training can further enhance their performance. The performance gain increases with a larger task graph size. 111 The code and datasets are available at https://github.com/WxxShirley/GNN4TaskPlan
@@ 1 Introduction
LLM-based agents have recently emerged as a rapidly growing field of research and are considered a significant step towards artificial general intelligence (AGI) [60, 7]. These agents have achieved remarkable successes across a variety of domains, as evidenced by their ability to address complex AI challenges (e.g., HuggingGPT [45]), excel in gaming environments (e.g., Voyager [57]), and drive innovation in chemical research (e.g., [5]). Within this burgeoning field, task planning in language agents emerges as a critical area of study. It involves LLMs autonomously interpreting user instructions, breaking user’s instructions in natural language into concrete and solvable sub-tasks, and then fulfilling the user’s request by executing each sub-task [42, 45, 44]. For instance, in the case of HuggingGPT [45], task planning involves invoking expert AI models from the HuggingFace website to solve complex AI tasks beyond the capabilities of GPT alone.
@@
Given its practical significance, numerous algorithms have been proposed, with a major focus on prompt design [44, 45, 32, 29, 59, 48, 70, 4, 62]. This paper proposes to explore an orthogonal direction, i.e., graph-learning-based approaches. In task planning, solvable sub-tasks can be naturally represented as a task graph, wherein each node corresponds to a distinct sub-task, and each edge signifies the dependencies between these sub-tasks. The crux of task planning, therefore, involves selecting a connected path or subgraph to satisfy the user’s request, which is a decision-making problem on graphs. Adopting this framework, we analyze the task planning capabilities of LLMs, specifically within the context of HuggingGPT [45]. Our empirical investigation uncovers that a considerable portion of planning failures can be ascribed to the LLMs’ inefficacy in accurately discerning the structure of the task graph. This finding presents intriguing questions from both theoretical and empirical perspectives. Theoretically, it initiates a discussion on the inherent limitations of LLMs in processing task graphs. Empirically, it highlights the urgent need for developing effective and efficient strategies to mitigate this deficiency and improve task planning performance.
@@
For the theoretical question, we first investigate the expressiveness of Transformer architectures when applied to graph tasks with sequential graph input, such as edge list representations, which is the graph input format for task planning. Our initial hypothesis is that the format of sequential graph input might not align with the inductive bias inherent to graph structures, potentially reducing expressiveness. Contrary to this hypothesis, it is proved that by taking edge lists as the input, a constant-width Transformer can solve graph decision-making problems by simulating dynamic programming algorithms on edge lists. Nevertheless, we find that LLMs’ solutions lack invariance under graph isomorphism, an important property for graph decision-making problems. In addition, the expressiveness is weakened if the attention is sparse, which is typically observed in LLMs [65]. Beyond expressiveness, we also examine the influence of auto-regressive loss, demonstrating that it introduces spurious correlations that can be harmful to graph decision-making tasks. These insights expose the inherent limitations of LLMs in task planning and, more broadly, in graph-related problems (e.g., the challenges in [14, 58, 33]).
@@
2 Preliminaries
In this section, we introduce task planning in language agents and the current LLM-based solutions.
2.1 Task Planning in Language Agents
We start with the definition of task planning with a concrete example of HuggingGPT [45]. In task planning, there is a pool of pre-defined tasks. Task planning inputs include this task pool and a user request. The user request is expressed in natural language, which is ambiguous and could encompass multiple complex tasks. The output is a sequence of tasks and the order of their invocation to address the user’s request.
[…]
2.2 Current LLM-based Solution to Task Planning
The current solution of task planning is purely based on LLMs and involves two stages [44, 45]. The first stage involves request decomposition, where a user’s ambiguous request is broken down into concrete steps via LLMs. For instance, the request illustrated in Figure 1 is decomposed into the following steps: (1) analyze the pose of the boy; (2) take that pose and generate a new image; (3) generate the caption for newly generated image; (4) convert the generated text into audio. The second stage is task retrieval. For each decomposed step, LLMs are employed to retrieve an appropriate task from the task pool and execute them in sequence. For example, “Analyze the pose of the boy” corresponds to Pose detection. The output tasks should be (1) Pose detection; (2) Pose-to-Image; (3) Image-to-Text; (4) Text-to-Speech. Figure 6 illustrates this procedure.
Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT)
3 Graph Formulation and Insights
3.1 Graph Formulation of Task Planning
In this subsection, we formulate the task planning as a decision-making problem on the task graph. The task graph is a special kind of text-attributed graphs and we define it as $G=(V,E,T)$. Each node $v\in V$ represents a pre-defined task in the task pool, associated with a text $t_{v}\in T$ describing its function (e.g., “Translation. Translation is the task of converting text from one language to another.”). Each edge $(u,v)\in E$ indicates a dependency between tasks (e.g., the output format of task $u$ matches the input format of task $v$). Task planning is to select a path or connected sub-graph on the task graph.
Viewed from this angle, task planning bears resemblance to traditional decision-making problems on graphs, such as planning for the shortest path. Compared with traditional planning, task planning in language agents involves diverse and open-ended goals due to the varied users’ personal requests. For example, on platforms like HuggingFace, users’ intentions span across video, text, and image domains. On the contrary, classic planning has a fixed goal for a given domain, which is often explicitly expressed by mathematical formulas [54, 35].
(a) Performance and Hallucination ratios
(b) Hallucination ratios across datasets
@@ 3.1 Graph Formulation of Task Planning
Figure 2: Illustration of (a) LLMs’ planning performance and hallucination in HuggingGPT, and (b) hallucination in relation to task graph size.
3.2 Failures of LLMs in Planning: Empirical Findings
@@
With the task graph at hand, we diagnose LLMs in task planning in Figure 2. We adopt the experimental settings as outlined in the work of HuggingGPT [45], where the prompts are specifically optimized for task planning on HuggingFace. The evaluation metric calculates the F1 score to assess the accuracy of the tasks identified by LLMs against the ground-truth tasks. Additionally, we report two task-graph-related metrics: the node hallucination ratio and the edge hallucination ratio. These metrics measure the frequency of non-existent nodes (i.e., tasks) and edges (i.e., dependencies) outputted by LLMs, respectively, indicative of the models’ misinterpretation of the graph input.
@@
We further explore whether the incidence of hallucinations correlates with the number of sub-tasks. The HuggingGPT dataset contains 23 sub-tasks, and our analysis is expanded to incorporate the UltraTool dataset [19], which consists of 260 sub-tasks. Figure 2(b) illustrates that the hallucination ratio increases with a larger task graph size.
3.3 Failures of LLMs in Planning: Theoretical Insights
In this subsection, we provide theoretical insights into the limitations of LLMs in processing task graphs. In contrast to previous graph learning approaches for graph decision-making problems, LLMs process the graph input by flattening it into a sequence and are trained using an auto-regressive loss. We will then examine the impact of these two factors.
How does sequential graph input impact the expressiveness? We consider general graph decision-making problems that can be resolved using dynamic programming (DP) as described in (2). The input comprises the edge list and initial states:
@@ 3.3 Failures of LLMs in Planning: Theoretical Insights
The intended output format is $u_{1}\ \text{Answer}[k][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[k][u_{n}]$. In existing studies, task graphs are often presented in (1), where the edge list is described by natural language and the initial states are the task descriptions of each task node, detailed in Appendix A.9 of [44].
As discussed in the previous subsection, task planning is a decision-making problem on the task graph. The decision-making problems on graphs are often solved by DP [3] and its general formulation is given by
$\text{Answer}[k][i]=f\left(\square_{j\in\mathcal{T}(i)}g(\text{Answer}[k-1][j],c[i][j])\right),$ (2)
where $\text{Answer}[k][i]$ is the solution to state $i$ in the $k$-th iteration, $c[i][j]$ is a cost associated with state $i$ and $j$, $\mathcal{T}(i)$ is the set of state can be transited to $i$, $\square_{j\in\mathcal{T}(i)}$ is an aggregation function such as MAX or SUM, and $f,g$ are task-specific update functions. We give the formulation of typical DP algorithms in Appendix D.1 including some NPC problems. For the decision-making problems on the text-attributed graphs (e.g., task planning), one may conceptualize them as DP in the feature space, as discussed in [67].
To our surprise, although the edge list input does not directly reflect the geometric structures of graphs, it enables Transformers to simulate DP efficiently, in terms of expressiveness, as demonstrated by the following theorem.
@@
(LLMs have enough expressiveness) Assume the input format is given in (1) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2 in Appendix. There exists a log-precision constant-depth and constant-width Transformer that simulates one step of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
The proof is presented in Appendix D.2. However, certain aspects of the proof’s constructions are challenging to be realized in Transformers that have been pretrained on natural language. First, the embedding process must be carefully filtered to ensure invariance under graph isomorphism. This invariance property does not align with the inductive biases inherent in natural language, making it difficult to achieve. Consequently, if an LLM can accurately produce the correct answer for a specific ordering of nodes, it might not maintain this accuracy after the nodes have been reordered (experiments given in Appendix D.3). Second, each token needs to synchronize its hidden states with all other tokens sharing the same token ID, which is of order $O(|V|)$. In practice, the attention trained from natural language is typically sparse [65], leading to intractability issues. The formal lower bound is provided in the following proposition and the proof is given in Appendix D.4.
@@
(Inductive bias of language hinders expressiveness) Assume the input format is described (1) and that the attention mechanism is limited to attending to a constant number of tokens. There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
How does auto-regressive Loss impact the generalization? Our investigation next focuses on the auto-regressive loss and considers the following scenario: given a fixed task graph, user data is collected to perform instruction tuning with next-token-prediction loss. For a tractable theoretical study, we conceptualize this issue as a path-planning problem, since the output of task planning is essentially a path. We consider the training dataset comprises input sequences of the form $s\ t\ s\ v_{1}\ v_{2}\ \cdots\ t$, where $s$ represents the source node, $t$ the target node, and the sequence $s\ v_{1}\ v_{2}\ \cdots\ t$ is a path that adheres to specified constraints. During testing, given the initial and target nodes $s$ and $t$, the model is expected to generate a path with the same constraint. Our findings indicate that auto-regressive loss can lead to the emergence of a frequency-based spurious correlation, as substantiated by the following theorem and the proof is given in Appendix D.5.
@@
To a human, finding a path from $a$ to $d$ simply involves concatenating the paths $a\ b\ c$ and $b\ c\ d$. However, auto-regressive loss fails under such circumstances. In task planning datasets, we indeed observe that the performance of fine-tuned LLMs is inferior to that of GNNs trained on the same dataset, as shown in Figure 3(b).
4 Integrating GNNs and LLMs for Planning
4.1 Motivations
In the last section, we find that a considerable portion of planning failures can be ascribed to the LLMs’ inefficacy in accurately discerning the structure of the task graph, due to the hallucination, the inductive bias of the attention, and next-token prediction loss. In contrast to LLMs, GNNs can strictly operate on the task graph, thereby avoiding hallucinations. Additionally, they leverage the graph structure as input, rather than flattening the graph into a sequence, thus overcoming the theoretical limitations discussed previously. Furthermore, GNNs have demonstrated proficiency in handling graph decision-making problems, both theoretically and empirically [67, 11, 23]. As a result, the simplest fix is to integrate GNNs into the task-planning algorithm.
In the following subsections, we propose both training-free and training-based approaches to enhance performance. Training-free methods are necessary when the available tasks are continuously changing, or new tasks are emerging constantly. This scenario is common when the task planning module is deployed in a new system. Once the task planning module has been deployed for a period, it becomes possible to collect users’ requests and label a small proportion of the data, enabling lightweight training-based methods.